OpenStudy (anonymous):

what is the ratio of 20 weeks : (to) 14 days

OpenStudy (anonymous):

20weeks=7*20 days so required ratio=7*20:14=10:1

OpenStudy (anonymous):

@saajidh

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

lol your welcome